# Invisible Labor in Open Source Software Ecosystems

### v1

@@ 4.4. Visibility Anchoring May Shift Beliefs About the Importance of Credit
Finally, we examine relationships between the quantitative variables of our survey. Calculating Pearson correlation coefficients between these variables reveals several trends (Fig. 5). Unsurprisingly, the likelihoods of receiving credit for specific tasks and from specific projects are highly positively correlated ($r=0.87$), as were participants’ satisfactions with the credit they receive ($r=0.73$, compare Fig. 3c & d). Credit frequency and participant satisfaction tended to correlate with one another (A, B, E, & F) and with full visibility (D). Partial visibility (C) also correlates with the former but less so, while negatively correlating with anchoring order (H, $r=-0.30$) and being not visible (I, $r=-0.27$). Indeed, negative correlations proved common for nonvisibility (A-F) and only positively correlated with anchoring to low visibility (H, $r=0.30)$. Anchoring to low visibility also correlated with importance of receiving credit (G, $r=0.26$), which in turn correlated with both credit frequencies ($r=0.23$ $r=0.20$, respectively).
@@
The trend between visibility and credit becomes more important in the context of one last observation. Recall that the importance of receiving credit (Fig. 5, G) correlated with anchoring order (H). The importance of receiving credit (Q16) appeared after the visibility questions (Q12-14) meaning that anchoring order had the potential to yield causal effects for subsequent questions. Fig. 6b shows that anchoring had a statistically significant effect on how important credit was to participants. On average, participants anchored to low visibility thought receiving credit was $30.5\%$ more important than participants anchored to high visibility ($95\%$ confidence interval of ±$\pm 19.5\%$, $p<0.002$, Adj. $R^{2}=0.061$ with robust standard errors). We discuss the implications of this and our other findings next.
@@ 5. Discussion
5.1. Findings & Implications
Labor becomes invisible labor when it is undercompensated, not fully visible, or both. Through a labor-diverse sample of OSS ecosystems, we found that invisible labor seems to pervade OSS ecosystems (Fig. 1b). Participants reported receiving credit for roughly half of their work. While this finding cannot speak to other forms of compensation (e.g. pay, social status, opportunities), credit is one of the least financially and socially expensive forms of compensation suggesting that this compensation rate is potentially an upper bound on compensation frequency.
Participants reported feeling only slightly satisfied with how often they receive credit and the mediums through which they receive it. This slight satisfaction suggests a possible explanation for why many individuals do not seek more credit. Many who were less satisfied with how much credit they receive listed sources of credit similar to those who were satisfied. However, satisfied individuals cited several sources that dissatisfied individuals cited less often including blogs (47.9% for satisfied, 5.3% for dissatisfied), presentations (37.5% vs. 10.5%), and project documentation (64.6% vs. 42.1%). It is possible, then, that these mediums might increase the efficacy of sharing information or the accessibility of information about OSS labor.
We then found that visibility anchoring shapes perceptions of the visibility of a person’s work. Asking participants about highly visible work (seen by $\geq 2$ people) first in our series of questions increased participant frequency perceptions of partially visible work (seen by 1 other person) and decreased frequency perceptions of low visibility work (seen by nobody else) compared to those whom we asked about low visibility first.
Our finding that visibility anchoring shifts labor visibility perceptions has two implications for OSS ecosystems. OSS communities tend to presume that work is visible because “openness means visibility” so to speak. This is not necessarily the case because visibility requires information to exist, get shared, and be accessible. As we have seen, OSS labor is only partially visible on average and roughly half does not receive credit. So presuming that openness creates visibility may actually bias OSS participants into perceiving invisible labor as less common than it is, thereby decreasing compensation efforts. The finding that visibility anchoring reduces credit importance perceptions further supports this likelihood. In addition, participants generally viewed partial visibility (seen by 1 other person) as more common than full visibility or nonvisibility. This confirms the precariousness of OSS labor noted by others (Eghbal, 2020) — here in the form of OSS labor visibility — for two reasons. First, if the singular external viewer of a task loses visibility into that task — by promotion, leaving the community, or even just forgetting — that labor becomes unknown to others. Second, some labor is already seen by nobody other than the worker. Either way, these tasks are invisible and therefore difficult for others to understand, compensate, and manage. Assuming this work is valuable to an OSS ecosystem, it reinforces the importance of accessible recordkeeping.
The correlation and regression results also imply that making OSS labor more visible probably requires giving individuals credit more often and more consistently. Yet even those who always receive credit did not report their work being completely visible. So though not all work should be fully visible, we join others who take the feminist position that all work should receive compensation (D’Ignazio and Klein, 2020). Invisible work — completely unknown to others — cannot receive compensation from anyone other than the individual performing the labor. If individuals have sufficient means (e.g. wealth, social status), they may be able to compensate themselves for performing invisible tasks, but most people cannot. We see this already in OSS when privileged individuals receive credit for activities that surpass community-defined thresholds of social status or centrality while less privileged individuals do not receive such credit (D’Ignazio and Klein, 2020; Young et al., 2021).
To remedy all of this, OSS ecosystems likely need more robust qualitative and quantitative documentation so that labor data exists; more effective dissemination practices so labor data gets shared; more easily understood data communication so labor data becomes accessible; and more credit in valued forms so all labor receives compensation. Treating all work as contributions may be an effective way to start (Casari et al., 2021; Allen et al., 2014).

### v2

@@ 4.4. Visibility Anchoring May Shift Beliefs About the Importance of Credit
Next, we examine relationships between the quantitative variables of our survey. Calculating Pearson correlation coefficients between these variables reveals several trends (Fig. 5). Unsurprisingly, the likelihoods of receiving credit for specific tasks and from specific projects are highly positively correlated ($r=0.87$), as were participants’ satisfactions with the credit they receive ($r=0.73$, compare Fig. 3c & d). Credit frequency and participant satisfaction tended to correlate with one another (A, B, E, & F) and with full visibility (D). Partial visibility (C) also correlates with the former but less so, while negatively correlating with anchoring order (H, $r=-0.30$) and being not visible (I, $r=-0.27$). Indeed, negative correlations proved common for nonvisibility (A-F) and only positively correlated with anchoring to low visibility (H, $r=0.30)$. Anchoring to low visibility also correlated with importance of receiving credit (G, $r=0.26$), which in turn correlated with both credit frequencies ($r=0.23$ $r=0.20$, respectively).
@@
The trend between visibility and credit becomes more important in the context of one last observation. Recall that the importance of receiving credit (Fig. 5, G) correlated with anchoring order (H). The importance of receiving credit (Q16) appeared after the visibility questions (Q12-14) meaning that anchoring order had the potential to yield causal effects for subsequent questions. Fig. 6b shows that anchoring had a statistically significant effect on how important credit was to participants. On average, participants anchored to low visibility thought receiving credit was $30.5\%$ more important than participants anchored to high visibility ($95\%$ confidence interval of ±$\pm 19.5\%$, $p<0.002$, Adj. $R^{2}=0.061$ with robust standard errors).
@@
To summarize, we found that labor is only partially visible and compensated on average (each about 50%). Further, we found that visibility anchoring shapes perceptions of the visibility of a person’s work. Asking participants about highly visible work (seen by $\geq 2$ people) first in our series of questions increased participant frequency perceptions of partially visible work (seen by 1 other person) and decreased frequency perceptions of low visibility work (seen by nobody else) compared to those whom we asked about low visibility first.
5. Qualitative Explanation of the Commonness of Invisible Labor
Given this quantitative evidence, we now consider factors that affect how commonly labor invisibility occurs (RQ2). Our free response questions asked participants about their experiences with how projects gave credit (Q17-19). In what follows, we analyze responses to these questions to identify factors that affect the prevalence of invisible labor, treating expressed experiences as expertise (Bardzell, 2010). We first corroborate our quantitative findings, providing qualitative evidence of labor invisibility from our sample. Then, our thematic analysis identifies a theme of attributing at cross-purposes — pursuing diverse motivations through conflicting attribution practices with limited resources — and its influence on the pervasiveness of invisible labor.
5.1. Experiences of Invisibility in OSS Ecosystems
Generally, participants written responses expressed dissatisfaction with the extents of visibility and compensation in OSS ecosystems, the constituents of invisible labor. A few respondents dissented from this view, though, expressing either apathy or hostility toward the idea of receiving credit. We address each respectively.
5.1.1. Visibility
Participants often recounted the non- and partial visibility of OSS labor. For example, participant P110 described how “contributions are seen by a handful of people at most which makes contributions almost invisible.” Similarly, P002 described how the work they do “only shows in project-specific mediums via being parsed out of commit messages, [which] doesn’t help with other things people might care about such as GitHub graphs.”
This second response exemplifies how technologies both enable and constrain visibility. Here, P002 implies that “GitHub graphs” achieve better visibility than “commit messages.” Interpreted through the elements of visibility (being recorded, shared, and accessible; see Sec. 2.1), P002 perceives a labor visualization tool (the GitHub graph) as increasing labor data accessibility for external stakeholders in a way that parsing one’s name from commit messages does not. This matches our understanding that recordkeeping alone, without sharing and accessibility of labor information, does not yield full visibility. That said, other participants took issue with the counting function performed by visibility tools like the GitHub graph:
Counting tasks or git issues can be super misleading. People can comment on an Issue and get credit just for participating, which should be appreciated. But there is no good solution for people who spend several hours writing an Issue or an Issue comment. Time spent on a task is not considered, nor are skill levels. Often programming/coding is seen as the only valid contribution type, and everything else is seen as trimmings and less valuable. (P104)
Another participant echoed these sentiments: “github showcases contributors to a repository in the sidebar, but only commits count as contributions, whereas activities like issue reporting, issue triaging, code review, translations performed in third-party platforms, spreading the word, etc., tend to be overlooked” (P82). [^9] GitHub has gradually expanded the activities that count toward the contribution graph to include issues, pull requests, discussions, and some commits. But we do not have evidence of the exact dates of when these changes were introduced, and to date it cannot account for activity outside the platform.
These quotes demonstrate how some OSS labor gets “overlooked” (i.e. recorded and shared less often), consistent with the literature (Geiger et al., 2021). Activities such as “issue reporting, issue triaging, code review, translations …spreading the word, etc.” are only “seen by a handful of people at most.” This happens partially because certain activities are “seen as trimmings” (i.e. not as real work) and as “less valuable.” Additionally, technologies make certain “valid contribution[s]” (especially “programming/coding”) more accessible by “showcas[ing]” them, although technologies also misrepresent certain dimensions of even legitimated activities (time, skill). Collectively, these choices have the effect of “misleading” people (both internal and external) about labor that takes place in an ecosystem. Put another way, decisions about what we make visible via recordkeeping, sharing, and accessibility often misrepresent the full extent of what gets done.
5.1.2. Compensation
Many participants also felt that OSS work received insufficient compensation (mostly referring to credit, but not only). Several participants described not receiving compensation on the basis of an individual’s status, such as community tenure status or residency status. On community tenure, for example, P024 responded with frustration that a project “didn’t [give me credit] at all. I made my first contribution to a browser and I was hoping to be listed anywhere, even if buried in some page of the browser’s system that no one ever finds organically.” By implication, they were not listed in those buried pages. On residency status, P009 described an experience common to many in minoritized regions who participate in global ecosystems: “Some people are less likely to receive credit based on timezone/location. For example, when Americans are giving out special credit, they tend to forget about Asians because they don’t see them in meetings a lot. Also, company loyalty messed up crediting people on multiple occasions.”
Others described their compensation experiences in starker terms: “It’s not just a credit/no credit issue, but a credit/no credit/erasure issue” (P131, emphasis added). While “erasure” is a visibility metaphor, the juxtaposition of “erasure” with “credit/no credit” here implies that erasure refers to taking away something owed, something compensatory that participants deserved. Depictions of “erasure” were common. One participant described how “maintainers would take credit for things they hadn’t done, and refrain to give any but bare credits and acknowledgement for other people’s work” (P129, emphasis added). Another described how “after 18 months of work on a project, nobody from the project leadership gave me any attribution or thanks” (P027). A third described how “people take the work and before it has your name on it, do some tangential re-spin, and commit in their name. What annoys me isn’t really the credit, but they act like a bad client, who isn’t paying you” (P062). A less generous interpretation might be one of theft. Another participant made this explicit through a metaphor of “slavery”:
Credits are good. But at some point you realize that while you’re being credited and such, you live with parents, and barely make \$150 a month to compensate them your living costs. Because by contributing to OSS and getting credit you’re in a reward loop that makes all proprietary work look like slavery under a mighty boss. (P101)
Participants described insufficient compensation in terms of harm. In their words, it “hurts people when it’s inconsistent, and not everyone is recognized in the same way” (P134). Some harms compare to affective pressures found in many workplaces: “In some of these scenarios where your role is not just creating a PR [Pull Request] with code contributions, I sometimes felt a pressure to keep up and make sure my role was still useful to the team” (P099). Other times, these pressures may have devolved into harassment: “Contributing contentious kind of work (design and Product management) is hard and very hostile at worst” (P096).
Beyond individual harms, giving insufficient compensation limits governing organizations, too. Labor retention grows from both giving credit (“Giving credit is important. Not giving credit can drive people away quickly,” P086) and giving it fairly (“Too many projects give credit to a small group [of workers], making it hard for the project to bring in new people,” P018). Some ecosystem leaders took this responsibility quite seriously, describing how they “deliberately try to acknowledge the work of others even if it downplays my own efforts. People need thanks. I used to send Christmas cards to the top contributors or new ones that needed a thanks” (P086). In sum, many participants felt insufficiently compensated and hurt as a result.
5.1.3. Dissenting Perspectives
Not all participants cared about labor invisibility, though. Some expressed relative indifference. They mentioned that “receiving credit for contributions to open source projects is certainly nice and makes me feel good, but it’s lower on the list of reasons why I contribute” (P040); that “this isn’t really something that concerns me..” (P045); and that “I do it because I want it, do not need credit” (P116). Others were more blunt:
If you come across anyone contributing to open source where recognition is important, they are the problem with open source. I contribute because I want to, because I need the feature, or because I want to see if I can do something. That’s how OS started. I stop contributing when I want to for my own reasons, not because I’m not getting enough street cred or recognition. (P020)
While a minority perspective, this participant was not alone in dissenting on the basis of personal enjoyment, learning, and own-use motivations.
In short, most participants described portions of their work as invisible, either for being insufficiently visible or insufficiently compensated. Not all care (tying into motivations), but most do. These descriptions provide a backdrop for our thematic analysis, which we turn to next.
5.2. Attributing at Cross-Purposes
Shaped by our survey questions, our thematic analysis identified one factor in our dataset that appears to influence the prevalence of invisible labor in OSS ecosystems — attributing at cross-purposes — with three subthemes related to participant motivations and attribution practices (Fig. 7). [^10] Other factors surely also influence labor invisibility. Here, we report only factors that we had evidence for as viewed through the theoretic lenses of this work. We discuss other possibilities in the Background (Sec. 2.3.3) and the Limitations (Sec. 6.2).
Again, motivations for participating in OSS vary and often coincide with one another (see Sec. 2.3.1). Our participants’ motivations were no exception. Affinity for community (i.e. kinship), career advancement, and pay were most prominent in our responses, though we also found evidence of interest in helping others (altruism) and improving tools (own-use). These motivations were often latent within responses about credit qualities that participants liked or disliked. Consider an excerpt from P062: “I guess it’s nice to have visible [data] to show when looking for jobs or in a negative phase looking for things to be positive about.” This excerpt contains an extrinsic career motive (“when looking for jobs”) tied to accessible labor information (“have visible [data] to show”), and a further unspecified motive (“looking for things to be positive about”). This latter motive is likely either intrinsic (e.g. altruism, kinship, fun) or internalized extrinsic (e.g. learning) because such motives tend toward greater affect (von Krogh et al., 2012) (as evidenced by the words “be positive about”).
While not mutually exclusive, individuals’ actions in pursuit of these motivations were often at cross-purposes with one another, though. By this, we mean that individuals and organizations might earnestly pursue several motivations simultaneously through attribution designs, but that such pursuits often got “complicated” (P004 & P058, separately) because efforts to advance one motivation frequently came at the expense of effectively pursuing another. This tension was particularly evident between three distinct forms of attribution (our subthemes) that participants spoke of: expressive attribution, instrumental attribution, and non-attribution. [^11] We draw our terminology for attribution from Balkundi & Harrison’s work on social network tie content (Balkundi and Harrison, 2006), though we acknowledge that other terminologies exist within CSCW to describe related concepts. Specifically with respect to Monroy-Hernández & colleagues (e.g. Monroy-Hernández et al., 2011), our instrumental attribution subtheme partially overlaps their “attribution” category, and our expressive attribution subtheme closely aligns with their “credit” category. We do not see a parallel to the non-attribution subtheme.
The first type, expressive attribution, describes genuine, personalized thanks for actions performed. This form of attribution is fundamentally relational in that individuals seek and appreciate sincere, person-to-person, non-automated expressions of affective states about labor performed. Those who sought it wanted not just “proper attribution of what was performed, [but also] why it was helpful” (P102). Expressive attribution likely provides individuals with “personal validation” (P045) of successfully fostering kinship, a fundamentally relational motive. At the same time, terms like “proper attribution” demonstrate how they did not view the giving of personalized thank you messages as distinct from internalized or extrinsic motivations. Sometimes participants sought public announcement of personalized statements (reputational motivation), for instance, which might yield invitations to be a “guest on a podcast, co-present at a conference, write a blog post” (P100). But in general, expressive attribution did not inherently require visibility so much as genuine thanks.
Next, instrumental attribution strives to systematically associate labor data with individuals who performed specific actions, and to make the data visible to those who might benefit from it. In contrast to the relational objectives of expressive attribution, this type of attribution is bureaucratic — not in the popular sense of needless process, so to speak, but rather because it seeks effective processes that systematically and transparently attribute labor, regardless of one’s relationship with others. Here, people often wanted attribution to be given “in an automated or semi-automated fashion” (P082) such that “everybody got credited, regardless of task/relevance/work, etc.” (P044) in a “specific and accurate” way (P042). Instrumental attribution seems to support altruism (i.e. “everybody got credited”), ideology (i.e. that work should be open), own-use (organizational control), and extrinsic (career, pay) motivations. It does this by systematizing recordkeeping, information sharing, and labor data accessibility in support of varied participant motivations.
Third is non-attribution, a type which was less common in our sample. Nevertheless, a few participants noted that they do not need (or even want) attribution. These individuals described their participation as “I do it because I want it [the software], do not need credit… I do it because I like it and not because others need to like it” (P116). These individuals see attribution as superfluous, unnecessary, or at least a lower priority. They self-describe as being motivated by own-use and fun. While labor visibility could hypothetically be an activity individuals associate with these motivations, we did not see evidence of it in our sample.
Participants expressed frustration at not receiving attribution in the way they wanted (i.e. the attribution type that best supports their motivations). Those seeking the genuine thanks of expressive attribution disliked the automated systems of instrumental attribution because they felt disingenuous (“anything automated feels fake,” P088), insufficiently personalized (“credits just don’t help to see how much damn effort it takes,” P101), and often produce inaccessible labor data (“Change log [attribution] wasn’t really useful,” P079). Those interested in instrumental attribution worried about how non-attribution hid work (“there was no way for others to see what had been done or not,” P057); for expressive attribution, they disliked the inconsistency (“projects give credit by accident, if at all,” P069), favoritism (“tends to be a few who garner all of the credit,” P112), and labor intensiveness (“often it’s a manual process that’s not easily scalable,” P082). Advocates of non-attribution expressed frustration with any attention to attribution (“anyone contributing to open source [for whom] recognition is important, they are the problem,” P020).
To reconcile this discord — particularly between proponents of expressive and instrumental attribution — many participants described mixed automated-and-manual attribution designs (“Mentioning in releasnotes [sic] and announcements on the blog,” P070). The objective therein was to both give personalized thanks to help people feel valued (kinship motivation) and thorough, accessible documentation that individuals could point to for career advancement. On the other hand, one person noted that their organization is discussing “whether individuals WANT to be named or credited — credit described in documentation offers that flexibility whereas automated systems would not account for this nuance and have the potential to extend beyond the comfort levels of participants” (P087). These balanced aspirations commonly gave way to pragmatic concerns, though:
Credit systems that give credit for all types of work, such as design, project management, event organizing, etc. [are ideal, but] too much manual effort required to give credit. Difficult to find mentors through the credit system. Hard to see who is doing what work. (P117)
The best situation is when the project is open to accept different methods of giving credit. For example, when in the past you needed 2 commits (in code) to get funds for a hackfest, now you can get funds even if you’re not a programmer and you contribute with marketing skills, community management etc. The most reliable metrics are stored in the memory of people who constantly hand out and know what’s happening. If they drop out, so does a lot of project value/knowledge/processes, and past credit they observed of other people. (P104)
From this, we see that limited individual resources (e.g. time, attention, knowledge, funds) and collective resources (like skilled members, accessible labor data, and capable technology) appear to constrain efforts to pursue everyone’s motivations to the fullest, here through attribution.
Altogether, this leaves OSS ecosystems in tension: While some participants do want greater visibility and compensation (instrumental attribution), others do not. Instead, they want to feel like part of a community (affirmed by expressive attribution) or to have fun doing whatever tasks they take pleasure in (non-attribution). Limited organizing, financial, and technical resources then mean that leaning into one type of attribution may help those who share similar motivations (as advanced by a particular form of attribution), but at the cost of pursuing or facilitating others. In this sense, efforts supporting one type of attribution (and the corresponding motivations) might not advance, or may even hinder, efforts to support other types of attribution (and their motivations). True, multiple attribution types can (and do) exist at once, but this too draws on limited resources, subsequently limiting abilities to pursue each form of attribution (and, once again, their motivations).
Expressive Attribution Instrumental Attribution Non-AttributionAB
Figure 7. Attribution types are often at cross-purposes with one another because pursuing motivations through one type of attribution (expressive, instrumental, or non-attribution) often constrains efforts to pursue other motivations. The triangle therefore symbolizes an “attribution space” in which one’s actions fall. For example, the more effort one puts toward expressive attribution (point A), the more difficult instrumental and non-attribution become. Likewise, equally supporting expressive and instrumental attribution (point B) for everyone would require significant effort and might frustrate those who prefer non-attribution.
[…]
The figure shows an equilateral triangle with the points of the triangle labeled “expressive attribution,” “instrumental attribution,” and “non-attribution,” the three subthemes of cross-purposed attribution.
Therefore, the confluence of diverse motivations and resource limitations frequently puts the attribution practices of individuals and organizations at odds with the motivations of others; that is, participants find their attribution practices at cross-purposes with one another. This includes efforts to reduce labor invisibility (primarily through instrumental attribution) which often limit efforts inspired by the motivations of other participants (as embodied in expressive and non-attribution). We represent this tension as a triangle delimiting the “attribution space” of possibilities (Fig. 7). This symbolically represents how efforts to pursue motivations through one or more attribution types (moving toward those points on the triangle) draw limited resources away from pursuing motivations through other attribution types. The triangle metaphor visualizes how the tension between attribution practices decreases labor visibility and compensation because many participants are motivated by interests that do not benefit from, and may even be hindered by, greater attention to labor visibility and compensation.
In this light, our finding for RQ1 that roughly half of OSS labor remains invisible becomes less surprising. When OSS managing organizations design attribution systems (however formally or informally), those systems differentially advance some motivations more than others (answering RQ2). Now we turn to the implications of these findings.
@@ 6. Discussion
As we have seen, labor becomes invisible labor when it is undercompensated or not fully visible (Fig. 1a). Through a labor-diverse sample of OSS ecosystems, we demonstrated that invisible labor makes up about half of the work taking place in OSS ecosystems (Fig. 1b) and that visibility anchoring shapes perceptions of the visibility of a person’s work. Resource limitations seem partially responsible for labor invisibility as evidenced by our exposé of the tension between attribution motivations (Fig. 7). We now discuss implications in the context of those who came before us, along with the work’s limitations and future work.
6.1. Implications
At some level, the commonness of invisibility in OSS is not surprising; researchers have recounted this since the inception of CSCW (Suchman, 1995; Star and Strauss, 1999). At the same time, showing that such a significant proportion of work is invisible, as we see elsewhere in CSCW (Toxtli et al., 2021), holds mixed potentials. Invisibility often shrouds valuable work expertise (Suchman, 1995; Star, 1991). So on one hand, our findings could increase the legitimacy of articulation work by corroborating the abundance of transferable knowledge embedded in OSS ecosystems (Jergensen et al., 2011). On the other hand, that 2 in 3 tasks are not fully visible reminds us that these critical ecosystems (National Science Foundation, 2024), and less powerful participants in them (c.f. Star, 1991; Star and Strauss, 1999), remain vulnerable. We found that work was most likely to be partially visible (seen by 1 other person), so if the singular external viewer of a task loses visibility into that task — by promotion, leaving the community (Jergensen et al., 2011), or even just forgetting — that labor becomes unknown to others. Second, some labor is already seen by nobody other than the worker as we showed. Either way, these tasks are invisible and therefore difficult for others to understand, compensate, and manage.
The abundance of studies that struggle with OSS labor and its invisibility, paired with our work, reminds us of a pragmatic consideration: that we cannot measure all of labor, certainly not to the desired level of quality. “No representation of the world is either complete or permanent” (Gerson and Star, 1986, p.257) and even the act of “making visible can incur invisibilities” (Star and Strauss, 1999, p.25). Even with improved qualitative and quantitative measures (as we too offer in this work), it may not be meaningful or desirable to document some tasks to ensure privacy, autonomy, and flexibility. Emotion regulation, for example, may not be helpful to make record of in some contexts (e.g. deliberation over design changes), though it certainly may be in others (e.g. social work, therapy, ecosystem moderation). [^12] Our thanks to an anonymous reviewer for motivating this point.
With an eye toward equitable, balanced, and secure specifications for “What should and shouldn’t be visible?” (Star and Strauss, 1999), responsive design strategies probably need participatory techniques for understanding ecosystems (Germonprez et al., 2017; Frey et al., 2019). A balanced design may therefore include participatory determination between representatives of an ecosystem’s different stakeholder groups over what is meaningful to make visible (tasks, durations, difficulties, expertises, authorship, etc.); how to make them visible (recordkeeping, sharing, and accessibility designs); what to make invisible; and how to keep it invisible. This would help ensure the visibility of mutually-valued information while balancing a group’s collective interests (motivations, security, privacy, information overload, etc.). Designing such a process for every ecosystem could be quite time consuming and ineffective at balancing stakeholder voices, though, so it may help for researchers to develop “playbooks” to ease (but not necessarily standardize) these efforts by helping stakeholders think through desired quantities and qualities of labor data to make visible and invisible (c.f. Google, 2024). This might be aided by designing sufficient flexibility into collaborative platforms such that specific ecosystems can create contextually-defined artifacts for monitoring qualities of interest (Meluso et al., 2022).
However, our finding on visibility anchoring implies that OSS participants may assume that work is more visible than it actually is because “openness seems like visibility” to many. While a common assumption, this is not the case because visibility requires information to be recorded, shared, and accessible (Stohl et al., 2016) while openness typically falls along dimensions of transparency-opacity (roughly synonymous with visibility), inclusion-exclusion, and concentrated-distributed decision rights (Splitter et al., 2023). So if participants presume that openness creates visibility, it may actually bias participants into believing that invisible labor is less common than it really is, thereby decreasing attention, legitimacy, and compensation efforts. Our finding that visibility anchoring reduced credit importance perceptions reinforces this likelihood. Instead, designing ecosystems around an assumption that most information is not visible — despite being labeled “open” — may be more prudent, particularly because it would incentivize stakeholders to think carefully about what they care enough about to make visible (and invisible) during participatory design processes.
Quantitatively and qualitatively, that an average 50% of OSS labor gets compensated holds significant implications, too. Specifically, reducing invisibility requires greater compensation (in credit and pay). Again, while not all work should be fully visible, we join others who take the feminist position that workers should receive appropriate compensation for their work (D’Ignazio and Klein, 2020), if an individual is so motivated (whether extrinsic, intrinsic, or internalized). If individuals have sufficient means (e.g. wealth, social status), they may be able to compensate themselves for performing invisible tasks, but most people cannot. We see this already in OSS when privileged individuals receive credit for activities that surpass community-defined thresholds of social status or centrality while less privileged individuals do not receive such credit (D’Ignazio and Klein, 2020; Young et al., 2021). But OSS is too important to society to ignore that participants are regularly undercompensated, even in expensive forms (like attribution). Per our qualitative results, then, compensation may involve some combination of expressive attribution, instrumental attribution, non-attribution, pay, opportunities, etc. How individuals are compensated might also benefit from participatory design, this time in the context of an ecosystem’s pragmatic constraints (e.g. individuals’ evolving time commitments, individual and collective funding availability). Together, these efforts may help increase participant satisfaction with credit mediums (or attribution types) and frequencies by addressing participant motivations and airing to more stakeholders the pragmatic constraints on the ecosystem.